OpenStudy (annipuppi):

determine whether the vectors in each pair are normal u=(2,2) and v=(4,-1) p=(3,-2) and q=(4,6) a=(7,-1) b=(1,7)

ganeshie8 (ganeshie8):

Hint : dot product

OpenStudy (annipuppi):

We only had one lesson over vectors so I don't really understand them at all...

OpenStudy (annipuppi):

@ganeshie8

ganeshie8 (ganeshie8):

Okay, what do you know about dot product ?

OpenStudy (annipuppi):

Pretty much nothing...

ganeshie8 (ganeshie8):

\(\color{red}{\vec{a} = (a_1, a_2)}\) \(\vec{b} = (b_1, b_2)\) \[\color{Red}{\vec{a}} \cdot \vec{b} = \color{red}{a_1}b_1 + \color{red}{a_2}b_2 \]

ganeshie8 (ganeshie8):

dot product is one way of defining multiplication of two vectors

ganeshie8 (ganeshie8):

to find the dot product of two vectors you just `multiply the corresponding components` and `add` them

OpenStudy (annipuppi):

Okay how do I know if the vectors are normal?

ganeshie8 (ganeshie8):

if the dot product of two vectors is 0, then the vectors are normal

ganeshie8 (ganeshie8):

simply find the dot product of given vectors and see if you get 0

OpenStudy (annipuppi):

so they aren't normal? I got -16

ganeshie8 (ganeshie8):

for what vectors ?

ganeshie8 (ganeshie8):

Lets see if these vectors are normal : \(\vec{u}=(2,2)\) \(\vec{v}=(4,-1)\) \(\vec{u}\cdot \vec{v} = 2(4) + 2(-1) = 8-2 = 6 \ne 0\) so the vectors \(\vec{u}\) and \(\vec{v}\) are not normal